In a basic Keynesian macroeconomic model, it is assumed that
Y = C + I
Where;
I = 250 and C = 0.75Y.
What is the equilibrium level of Y? What increase in I would be needed to cause Y to increase to 1,200?
1
Expert's answer
2019-09-30T09:41:17-0400

We know that Y = C + I

where I = 250 and C = 0.75Y


"\therefore Y = 0.75Y + 250"


"\therefore Y - 0.75Y = 250"


"\therefore 0.25Y = 250"


"\therefore Y = \frac{250}{0.25}"

Thus, Y = 1000 (Equilibrium level of Y)


In order to assess the increase needed in I to cause Y to increase to 1200, we can use the following methodology:

Let a be the increase needed in I to cause Y to increase to 1200.


Thus the updated values for the above equation of Y = C + I are:

Y = 1200,

C = 0.75Y = 0.75* 1200 = 900,

I = 250 + a


Y = C + I implies:

1200 = 900 + (250 + a)

1200 = 1150 + a

1200 - 1150 = a


"\therefore a = 50"


There should be an increase of 50 in I to cause Y to increase to 1200.